OpenStudy (anonymous):

What is cos of 45 degrees?

OpenStudy (gorv):

it is standard value

OpenStudy (gorv):

\[\cos 45=\frac{ 1 }{ \sqrt 2 }\]

OpenStudy (anonymous):

Thank you!!